OpenStudy (susanna):

Can someone check my answer? Find cos x if sin x cot x = 4 I think cos x = 4

OpenStudy (susanna):

\[\sin x \frac{ \cos x }{ \sin x }=4\]

OpenStudy (susanna):

cos x = 4 Am I right?

OpenStudy (mehek.15):

correct

OpenStudy (susanna):

Thanx, it WAS correct!